An equipment has initial cost of $ 80,887 and is expected to have a $ 13,603 SV when replaced after 38,571 hours of use. Find its BV after 17,426 hours of operation using SL method.
Answer #1

Answer:

= Depreciation of an equipment

= ( Initial value - Salvage value ) * Used hours / Total Hours

= ( 80887 - 13603 ) * 17426 / 38571

= 67,284‬ * 0.45

= $30,398

Book Value = Initial Value - Total depreciation = 80887 - 30398 = $50,489‬

Thus book value is $ 50,489‬